OpenStudy (anonymous):

Find sin2θ, cos2θ, and tan2θ: csc=3/1 270<θ<360

OpenStudy (anonymous):

@calculusfunctions -- if you're free, I have a tough one haha

OpenStudy (calculusfunctions):

Given:\[\csc \theta =3\]where\[270°<\theta <360\]Find sin 2θ, cos 2θ, and tan 2θ. Is this the intended question?

OpenStudy (anonymous):

I believe it's 3/1 but my paper got smudged a bit so it's the best I have to go with.. so in short, yes.

OpenStudy (calculusfunctions):

3/1 = 3 Now first draw the diagram in the appropriate quadrant Now label the diagram and show me.